### v1

@@ D.2 Proving Theorem 9: Main Result for Without Replacement Orderings
$2\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})$ $=\mathop{\mathbb{E}}_{\tau}\left\|{\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({k}\right)}}\right)\mathbf{w}_{k-1}}\right\|^{2}\leq\left\|{\mathbf{w}_{\star}}\right\|^{2}\mathop{\mathbb{E}}_{\tau}a_{k}\leq\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\sum_{t=2}^{k}\mathop{\mathbb{E}}_{\tau}a_{t}$
$=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\sum_{t=2}^{k}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)-\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}{\mathbf{P}}_{\tau\left({1}\right)}{\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)}\right]$
$\left[\begin{subarray}{c}\text{exchangeable}\end{subarray}\right]$ $=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\sum_{t=2}^{k}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t-1}\right)}}\right)-\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)}\right]$
$\left[\begin{subarray}{c}\text{telescoping}\end{subarray}\right]$ $=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}}\right)-\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({k}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({k}\right)}}\right)}\right]$
$\leq\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}}\right)}\right]=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}\left({d-\bar{r}}\right)}{k-1}.$
For the second, parameter independent term, note that from Lemma 6, $f_{m}(\cdot)$ is $1$-smooth for all $m\in[T]$, and recall that the iterates $\mathbf{w}_{t}$ follow the SGD dynamics with $\eta=1$ (Reduction 2). Hence, by Lemma 33, without-replacement SGD with $\beta=\eta=1$ guarantees that after $k\geq 1$ gradient steps:
@@
$\mathop{\mathbb{E}}_{\tau}\left[{F_{\tau}\left({k}\right)}\right]$ $\leq 2R^{2}\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}\leq R^{2}\min\left({\frac{2e\left\|{\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k-1}},\,\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}\left({d-\bar{r}}\right)}{k-1}}\right)+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$
@@ Appendix E Proofs of Last-Iterate SGD Bounds (Section 5)
In this section we provide proofs and full technical details of our upper bounds for least squares SGD. We begin by recording a few elementary well-known facts, which can be found in e.g., Bubeck (2015). We provide proof for completeness.
Lemma 30 (Fundamental regret inequality for gradient descent).
@@
which completes the proof.
Lemma 31 (Descent lemma).
@@
The second claim now follows by using the fact that $f(\mathbf{w}_{\star})\leq f(\mathbf{w}^{+})$.
E.1 Proofs for Section 5.1
@@
  • Recall Lemma 11.Consider the $\beta$-smooth, realizable Setup 1, and let $T\geq 1$, $(i_{0},\ldots,i_{T})\in\mathcal{I}^{T+1}$ be an arbitrary sequence of indices in $\mathcal{I}$, and ${\mathbf{w}_{0}\in\mathbb{R}^{d}}$ be an arbitrary initialization. Then, the gradient descent iterates given by ${\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\eta\nabla f(\mathbf{w}_{t};i_{t})}$ for a step size $\eta<2/\beta$, hold:
@@
Denote $f_{t}(\mathbf{w})\triangleq f(\mathbf{w};i_{t})$, and observe by Lemma 30;
@@
where the second inequality follows from Lemma 31. On the other hand, by Lemma 32,
@@
and the result follows after dividing by $(2-\eta\beta)$.
  • Recall Lemma 12.Consider the $\beta$-smooth, realizable Setup 1. Let $T\geq 1$. Assume $\mathcal{P}$ is a distribution over $\mathcal{I}^{T+1}$ such that for every $0\leq t\leq\tau_{1}\leq\tau_{2}\leq T$, the following holds: For any ${i_{0},\ldots i_{t-1}\in\mathcal{I}^{t}},i\in\mathcal{I}$, $\Pr(i_{\tau_{1}}=i|i_{0},\ldots,i_{t-1})=\Pr(i_{\tau_{2}}=i|i_{0},\ldots,i_{t-1})$. Then, for any initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, with-replacement SGD (Eq. 2) with step-size $\eta<2/\beta$, holds:
@@
Denote $f_{t}(\mathbf{w})\!\triangleq\!f(\mathbf{w};i_{t}),\,\mathbf{g}_{t}\!\triangleq\!\nabla f_{t}(\mathbf{w}_{t})$, and observe that by Lemma 30, $\forall\mathbf{z}\in\mathbb{R}^{d},\,t\leq T$ (w.p. $1$):
@@
By Lemma 32, this implies for any $\gamma>0$:
$\sum_{t=T-k}^{T}(2-\gamma-\eta\beta)f_{t}(\mathbf{w}_{t})-\left(\frac{1}{\gamma}-\eta\beta\right)f_{t}(\mathbf{z})$ $=\sum_{t=T-k}^{T}\!\Big{(}{(2-\gamma)f_{t}(\mathbf{w}_{t})-\frac{1}{\gamma}f_{t}(\mathbf{z})}\Big{)}+\eta\beta\!\sum_{t=T-k}^{T}f_{t}(\mathbf{z})-f_{t}(\mathbf{w}_{t})$
[Lemma 32]delimited-[]Lemma $\end{subarray}\right]$ $\leq\sum_{t=T-k}^{T}\left\langle\mathbf{g}_{t},\mathbf{w}_{t}-\mathbf{z}\right\rangle+\eta\beta\sum_{t=T-k}^{T}f_{t}(\mathbf{z})-f_{t}(\mathbf{w}_{t})$
@@
and by the assumption on the distribution $\mathcal{P}$ it follows that $\mathbb{E}f_{T-k}(\mathbf{w}_{T-k})=\mathbb{E}f_{t}(\mathbf{w}_{T-k})$ for any $t\geq T-k$. Thus, combined with our previous display,
$\mathbb{E}S_{k}$ $\leq\frac{1}{\big{(}{k+1}\big{)}\gamma(2-\gamma-\eta\beta)}\sum_{t=T-k}^{T}\mathbb{E}f_{t}(\mathbf{w}_{T-k})=\frac{1}{\big{(}{k+1}\big{)}\gamma(2-\gamma-\eta\beta)}\sum_{t=T-k}^{T}\Big{(}{(k+1)\mathbb{E}S_{k}-k\mathbb{E}S_{k-1}}\Big{)}$
@@
E.2 Proving Theorem 13
  • Recall Theorem 13.Consider the $\beta$-smooth, realizable Setup 1. Define for all $T\!\geq\!2$, $\hat{f}_{0:T}(\mathbf{w})\!\triangleq\!\frac{1}{T+1}\sum_{t=0}^{T}f(\mathbf{w};\pi_{t})$. Then, without-replacement SGD (Eq. (3)) with step-size $\eta<{2}/{\beta}$, holds:
@@ E.2 Proving Theorem 13
    where $D\triangleq\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|$. In particular,
    $\eta=\frac{1}{\beta\log T}\Longrightarrow\mathbb{E}\hat{f}_{0:T-1}(\mathbf{w}_{T})\leq{\displaystyle\frac{14\beta D^{2}\log T}{T}}\,,\quad\quad\eta=\frac{1}{\beta}\Longrightarrow\mathbb{E}\hat{f}_{0:T-1}(\mathbf{w}_{T})\leq\frac{7\beta D^{2}}{\sqrt[4]{T}}\,.$
The proof of our theorem follows by a combination of two lemmas. The first, stated below, establishes a bound on the expected “next sample” loss and follows immediately by combining Lemmas 12 and 11 (notice that $\eta<\frac{2}{\beta}\Longrightarrow\exp\big{(}{\eta\beta\big{(}{1-\frac{\eta\beta}{4}}\big{)}}\big{)}\mapsto\exp\big{(}{z\big{(}{1-\frac{z}{4}}\big{)}}\big{)}$ for $z\in(0,2)$, which is monotonic increasing and upper bounded by $e$).
Lemma 33.
@@
The proof of Lemma 34 builds on an algorithmic stability argument similar to that given in Lei and Ying (2020), combined with the without-replacement stability framework proposed by Sherman et al. (2021). Before turning to the proof given in the next subsection, we quickly prove Theorem 13.
Proof (of Theorem 13).
By Lemmas 34 and 33,
$\mathbb{E}_{\pi}\hat{f}_{0:T}(\mathbf{w}_{T})$ $\leq 2\mathbb{E}_{\pi}f(\mathbf{w}_{T};\pi_{T})+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{T+1}\leq\frac{e\cdot\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\eta(2-\eta\beta)}T^{\eta\beta\left(1-\frac{\eta\beta}{4}\right)-1}+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{T+1}.$
@@
$\frac{eD^{2}T^{\eta\beta\left(1-\eta\beta/4\right)-1}}{\eta(2-\eta\beta)}=\frac{e\beta D^{2}\log T}{T(2-\frac{1}{\log T})}T^{\frac{1}{\log T}\left(1-\frac{1}{4\log T}\right)}=\frac{\beta D^{2}\log T}{T}\frac{\exp\left(2-\frac{1}{4\log T}\right)}{2-\frac{1}{\log T}}\leq\frac{10\beta D^{2}\log T}{T}.~{}~{}~{}~{}~{}~{}~{}$
E.2.1 Proving Lemma 34
@@
The proof follows from observing that the random variables $f(\mathbf{w}_{\tau};\pi_{i})$ and $f(\mathbf{w}_{\tau}^{(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{i})$ are distributed identically (the indices $\pi_{i},\pi_{t}$ are exhangeable). Formally, let $\Pi(\mathcal{I})\triangleq\left\{\pi\in\mathcal{I}\leftrightarrow\mathcal{I}\right\}$ be the set of all permutations over $\mathcal{I}$, and observe
@@
the result follows.
Our next lemma, originally given in Sherman et al. (2021, Lemma 2 therein), can be thought of as a without-replacement version of the well known stability $\iff$ generalization argument of the with-replacement sampling case (Shalev-Shwartz et al., 2010; Hardt et al., 2016).
Lemma 36.
@@
We have, by definition of $\hat{f}_{0:t-1}$ and Lemma 35:
$\mathbb{E}_{\pi}\left[\hat{f}_{0:t-1}(\mathbf{w}_{t})\right]=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t};\pi_{i})\right]=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t}^{(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{i})\right]=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t}^{(i\leftrightarrow t)};\pi_{t})\right],$
where the last equality is immediate since by definition, $\pi(i\leftrightarrow t)_{i}=\pi_{t}$. The claim now follows by linearity of expectation.
We are now ready to prove our main lemma. We note that the proof shares some features with that of the with-replacement case (Lemma 38).
Proof (of Lemma 34).
@@
Next, we bound $\left\|{\mathbf{w}_{t}^{(i\leftrightarrow t)}-\mathbf{w}_{t}}\right\|^{2}$. For any $10\leq\tau\leq t-$ 1, we denote $f_{\tau}\triangleq f(\cdot;\pi_{\tau})$, $f_{\tau}^{(i\leftrightarrow t)}\triangleq f(\cdot;\pi(i\leftrightarrow t)_{\tau})$. Observe that for any $\tau$ such that $\tau\neq i$, we have $f_{\tau}=f_{\tau}^{(i\leftrightarrow t)}$, thus, by the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ (see Lemma 3.6 in Hardt et al., 2016):
@@
and by Lemma 35 $\mathbb{E}f_{i}(\mathbf{w}_{i})=\mathbb{E}f_{i}^{(i\leftrightarrow t)}(\mathbf{w}_{i}^{(i\leftrightarrow t)})$. Hence,
@@
which, when combined with Proof yields:
@@
Finally, by the regret bound given in Lemma 11, $\sum_{i=0}^{t-1}f_{i}(\mathbf{w}_{i})\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta(2-\eta\beta)}$, and therefore,
@@
which completes the proof.
Appendix F Proofs of Extensions (Section 6)
  • Recall Reduction 3.Consider $T$ arbitrary (nonempty) closed convex sets $\mathcal{C}_{1},\dots,\mathcal{C}_{T}$, initial point $\mathbf{w}_{0}\in\mathbb{R}^{d}$, and ordering $\tau$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2},\forall m\in\left[{T}\right]$. Then,

### v2

@@ C.1 Key Properties and Auxiliary Lemmas
$2\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})=\mathop{\mathbb{E}}_{\tau}\left\|{\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({k}\right)}}\right)\mathbf{w}_{k-1}}\right\|^{2}\leq\left\|{\mathbf{w}_{\star}}\right\|^{2}\mathop{\mathbb{E}}_{\tau}a_{k}\leq\tfrac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\sum_{t=2}^{k}\mathop{\mathbb{E}}_{\tau}a_{t}$
$=\tfrac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\sum_{t=2}^{k}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)-\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)}\right]$
$\left[\begin{subarray}{c}\text{exchangeability}\end{subarray}\right]$ $=\tfrac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\sum_{t=2}^{k}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t-1}\right)}}\right)-\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)}\right]$
$\left[\begin{subarray}{c}\text{telescoping}\end{subarray}\right]$ $=\tfrac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}}\right)-\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({k}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({k}\right)}}\right)}\right]$
$\leq\tfrac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}}\right)}\right]=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}\left({d-\bar{r}}\right)}{k-1}.$
For the second, parameter independent term, note that from Lemma 3.1, $f_{m}(\cdot)$ is $1$-smooth for all $m\in[T]$, and recall that the iterates $\mathbf{w}_{t}$ follow the SGD dynamics with $\eta=1$ (Reduction 2). Hence, by Lemma E.5, without-replacement SGD with $\beta=\eta=1$ guarantees that after $k\geq 1$ gradient steps:
@@
$\mathop{\mathbb{E}}_{\tau}\left[{F_{\tau}\left({k}\right)}\right]$ $\leq 2R^{2}\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$
$\leq R^{2}\min\left({\frac{2e\left\|{\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k-1}},\,\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}\left({d-\bar{r}}\right)}{k-1}}\right)+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$
@@ Appendix E Proofs of Last-Iterate SGD Bounds (Section 5)
In this section we provide proofs and full technical details of our upper bounds for least squares SGD. We begin by recording a few elementary well-known facts, which can be found in e.g., Bubeck [8]. We provide proof for completeness.
Lemma E.1 (Fundamental regret inequality for gradient descent).
@@
    which completes the proof. $\blacksquare$
    Lemma E.2 (Descent lemma).
@@
    The second claim now follows by using the fact that $f(\mathbf{w}_{\star})\leq f(\mathbf{w}^{+})$. $\blacksquare$
E.1 Proofs for With Replacement Orderings
@@
    $\blacksquare$
  • Recall Lemma 5.2.Consider the $\beta$-smooth, realizable Setup 1, and let $T\geq 1$, $(i_{0},\ldots,i_{T})\in\mathcal{I}^{T+1}$ be an arbitrary sequence of indices in $\mathcal{I}$, and ${\mathbf{w}_{0}\in\mathbb{R}^{d}}$ be an arbitrary initialization. Then, the gradient descent iterates given by ${\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\eta\nabla f(\mathbf{w}_{t};i_{t})}$ for a step size $\eta<2/\beta$, hold:
@@
    Denote $f_{t}(\mathbf{w})\triangleq f(\mathbf{w};i_{t})$, and observe by Lemma E.1;
@@
    where the second inequality follows from Lemma E.2. On the other hand, by Lemma E.3,
@@
    and the result follows after dividing by $(2-\eta\beta)$. $\blacksquare$
  • Recall Lemma 5.3.Consider the $\beta$-smooth, realizable Setup 1. Let $T\geq 1$. Assume $\mathcal{P}$ is a distribution over $\mathcal{I}^{T+1}$ such that for every $0\leq t\leq\tau_{1}\leq\tau_{2}\leq T$, the following holds: For any ${i_{0},\ldots i_{t-1}\in\mathcal{I}^{t}},i\in\mathcal{I}$, $\Pr(i_{\tau_{1}}=i|i_{0},\ldots,i_{t-1})=\Pr(i_{\tau_{2}}=i|i_{0},\ldots,i_{t-1})$. Then, for any initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, with-replacement SGD (Eq. 2) with step-size $\eta<2/\beta$, holds:
@@
    Denote $f_{t}(\mathbf{w})\!\triangleq\!f(\mathbf{w};i_{t}),\,\mathbf{g}_{t}\!\triangleq\!\nabla f_{t}(\mathbf{w}_{t})$, and observe that by Lemma E.1, $\forall\mathbf{z}\in\mathbb{R}^{d},\,t\leq T$ (w.p. $1$):
@@
    By Lemma E.3, this implies for any $\gamma>0$:
    $\sum_{t=T-k}^{T}(2-\gamma-\eta\beta)f_{t}(\mathbf{w}_{t})-\left(\frac{1}{\gamma}-\eta\beta\right)f_{t}(\mathbf{z})$
    $=\sum_{t=T-k}^{T}\!\Big{(}{(2-\gamma)f_{t}(\mathbf{w}_{t})-\frac{1}{\gamma}f_{t}(\mathbf{z})}\Big{)}+\eta\beta\!\sum_{t=T-k}^{T}f_{t}(\mathbf{z})-f_{t}(\mathbf{w}_{t})$
    [Lemma E.3]delimited-[]Lemma $\end{subarray}\right]$ $\leq\sum_{t=T-k}^{T}\left\langle\mathbf{g}_{t},\mathbf{w}_{t}-\mathbf{z}\right\rangle+\eta\beta\sum_{t=T-k}^{T}f_{t}(\mathbf{z})-f_{t}(\mathbf{w}_{t})$
@@
    and by the assumption on the distribution $\mathcal{P}$ it follows that $\mathbb{E}f_{T-k}(\mathbf{w}_{T-k})=\mathbb{E}f_{t}(\mathbf{w}_{T-k})$ for any $t\geq T-k$.
    Thus, combined with our previous display,
    $\mathbb{E}S_{k}$ $\leq\frac{1}{\big{(}{k+1}\big{)}\gamma(2-\gamma-\eta\beta)}\sum_{t=T-k}^{T}\mathbb{E}f_{t}(\mathbf{w}_{T-k})$
    $=\frac{1}{\big{(}{k+1}\big{)}\gamma(2-\gamma-\eta\beta)}\sum_{t=T-k}^{T}\Big{(}{(k+1)\mathbb{E}S_{k}-k\mathbb{E}S_{k-1}}\Big{)}$
@@
    $\blacksquare$
E.2 Extending the SGD Bounds to Without Replacement Orderings
Here, we extend Theorem 5.1 to a without-replacement setting. Specifically, we consider gradient descent under a random permutation of the $T$ tasks. That is, for some initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, step size $\eta>0$, and $\pi_{t}\sim\textrm{Unif}(\mathcal{I})$ sampled without replacement,
Theorem E.4.
Last-Iterate Bound for Realizable Regression Without Replacement Consider the $\beta$-smooth, realizable Setup 1. Define for all $T\!\geq\!2$, $\hat{f}_{0:T}(\mathbf{w})\!\triangleq\!\frac{1}{T+1}\sum_{t=0}^{T}f(\mathbf{w};\pi_{t})$. Then, without-replacement SGD (Eq. (7)) with step-size $\eta<{2}/{\beta}$, holds:
@@ E.2 Extending the SGD Bounds to Without Replacement Orderings
where $D\triangleq\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|$. In particular, for $\eta=\frac{1}{\beta\log T}$ yields ${\frac{14\beta D^{2}\log T}{T}}$ and $\eta=\frac{1}{\beta}$ yields $\frac{7\beta D^{2}}{\sqrt[4]{T}}$.
The proof, given next, is based on the algorithmic stability of SGD [6, 61, 25], and more specifically, on a variant of stability, suitable for without replacement sampling [63, 35].
The proof of our theorem follows by a combination of two lemmas. The first, stated below, establishes a bound on the expected “next sample” loss and follows immediately by combining Lemmas 5.3 and 5.2 (notice that $\eta<\frac{2}{\beta}\Longrightarrow\exp\big{(}{\eta\beta\big{(}{1-\frac{\eta\beta}{4}}\big{)}}\big{)}\mapsto\exp\big{(}{z\big{(}{1-\frac{z}{4}}\big{)}}\big{)}$ for $z\in(0,2)$, which is monotonic increasing and upper bounded by $e$).
Lemma E.5.
@@
The proof of Lemma E.6 builds on an algorithmic stability argument similar to that given in Lei and Ying [38], combined with the without-replacement stability framework proposed by Sherman et al. [63]. Before turning to the proof given in the next subsection, we quickly prove Theorem E.4.
  • Proof of Theorem E.4.
    By Lemmas E.6 and E.5,
    $\mathbb{E}_{\pi}\hat{f}_{0:T}(\mathbf{w}_{T})$ $\leq 2\mathbb{E}_{\pi}f(\mathbf{w}_{T};\pi_{T})+\tfrac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{T+1}\leq\tfrac{e\cdot\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\eta(2-\eta\beta)}T^{\eta\beta\left(1-\frac{\eta\beta}{4}\right)-1}+\tfrac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{T+1}.$
@@
    $\tfrac{eD^{2}T^{\eta\beta\left(1-\eta\beta/4\right)-1}}{\eta(2-\eta\beta)}=\frac{e\beta D^{2}\log T}{T(2\!-\!\frac{1}{\log T})}T^{\frac{1}{\log T}\left(1-\frac{1}{4\log T}\right)}=\tfrac{\beta D^{2}\log T}{T}\frac{\exp\left(2-\frac{1}{4\log T}\right)}{2-\frac{1}{\log T}}\leq\frac{10\beta D^{2}\log T}{T}.$
    $\blacksquare$
E.2.1 Proving Lemma E.6
@@
    The proof follows from observing that the random variables ${f(\mathbf{w}_{\tau};\pi_{i})}$ and $f(\mathbf{w}_{\tau}^{(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{i})$ are distributed identically (the indices $\pi_{i},\pi_{t}$ are exchangeable). Formally, let $\Pi(\mathcal{I})\triangleq\left\{\pi\in\mathcal{I}\leftrightarrow\mathcal{I}\right\}$ be the set of all permutations over $\mathcal{I}$, and observe
@@
    the result follows. $\blacksquare$
Our next lemma, originally given in Sherman et al. [63, Lemma 2 therein], can be thought of as a without-replacement version of the well known stability $\iff$ generalization argument of the with-replacement sampling case [61, 25].
Lemma E.8.
@@
    We have, by definition of $\hat{f}_{0:t-1}$ and Lemma E.7:
    $\mathbb{E}_{\pi}\left[\hat{f}_{0:t-1}(\mathbf{w}_{t})\right]$ $=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t};\pi_{i})\right]$
    $=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t}^{(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{i})\right]=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t}^{(i\leftrightarrow t)};\pi_{t})\right],$
    where the last equality is immediate since by definition, $\pi(i\leftrightarrow t)_{i}=\pi_{t}$. The claim now follows by linearity of expectation. $\blacksquare$
We are now ready to prove our main lemma. We note that the proof shares some features with that of the with-replacement case (Lemma F.2).
  • Proof of Lemma E.6.
@@
    Next, we bound $\left\|{\mathbf{w}_{t}^{(i\leftrightarrow t)}-\mathbf{w}_{t}}\right\|^{2}$. For any $10\leq\tau\leq t-$ 1, we denote $f_{\tau}\triangleq f(\cdot;\pi_{\tau})$, and $f_{\tau}^{(i\leftrightarrow t)}\triangleq f(\cdot;\pi(i\leftrightarrow t)_{\tau})$. Observe that for any $\tau$ such that $\tau\neq i$, we have $f_{\tau}=f_{\tau}^{(i\leftrightarrow t)}$, thus, by the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ (see Lemma 3.6 in 25):
@@
    and by Lemma E.7 $\mathbb{E}f_{i}(\mathbf{w}_{i})=\mathbb{E}f_{i}^{(i\leftrightarrow t)}(\mathbf{w}_{i}^{(i\leftrightarrow t)})$. Hence,
@@
    which, when combined with Proof of Lemma E.6. yields:
@@
    Finally, by the regret bound given in Lemma 5.2, $\sum_{i=0}^{t-1}f_{i}(\mathbf{w}_{i})\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta(2-\eta\beta)}$, and therefore,
@@
    which completes the proof. $\blacksquare$
Appendix F Supplementary Material for the Extension Section (Section 6)
  • Recall Reduction 3.Consider $T$ arbitrary (nonempty) closed convex sets $\mathcal{C}_{1},\dots,\mathcal{C}_{T}$, initial point $\mathbf{w}_{0}\in\mathbb{R}^{d}$, and ordering $\tau$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})}\right\|^{2},\forall m\in\left[{T}\right]$. Then,